The Graduate Record Examination (GRE) is a standardized test that students usually take before entering graduate school. According to the document Interpreting Your GRE Scores, a publication of the Educational Testing Service, the scores on the verbal portion of the GRE are approximately normally distributed with a mean of 462 points and a standard deviation of 119 points. Find the 95 th percentile for this exam score. please show work.
Answer #1

Given that,

mean = = 462

standard deviation = = 119

Using standard normal table,

P(Z < z) =95 %

= P(Z < z) = 0.95  

= P(Z <1.64 ) = 0.95

z = 1.64 Using standard normal z table,

Using z-score formula  

x= z * +

x=1.64 *119+462

x= 657.16